### v1

# Imaging the Femtoscopic Freeze-out Source in Heavy Ion Collisions

Abstract
By combining the femtoscopic interferometry and the optical deblurring algorithm, we implement a novel method to image the source formed in HICs, while the interaction strength between the particle pair can be simultaneously determined. The spatial distribution of the emission source has been reconstructed for protons ($p$) and antiprotons ($\bar{p}$) from the respective $pp$ and $\bar{p}\bar{p}$ correlation functions in Au+Au collisions at $\sqrt{S_{\rm NN}}=200$ GeV. Within experimental uncertainties, protons and antiprotons share the same freeze-out distribution showing higher density in the center compared to the widely assumed Gaussian shape. The results evidence the matter-antimatter symmetry in coordinate space at the moment of freeze-out before the nucleons are fully randomized in the collisions.
It is accepted that matter and antimatter are created in equal amount in the first moments after the Big Bang. Yet today, we observe a universe filled with matter far exceeding antimatter. This puzzling phenomenon is related to the profound symmetry breaking in nature and remains a challenge in nuclear and particle physics. On modern accelerators, HICs create conditions akin to those in the early universe. So after the success of observing the production of antimatter [1, 2, 3], people can further investigate the spatial evolution and decoupling of the matter and antimatter on the femtoscopic scale.
The HBT intensity interferometry, invented by Hanbury-Brown and Twiss to detect the angular radius of Siris [4], has been extensively applied in nuclear physics [5, 6] to extract the temporal-spatial feature of the emission source formed in HICs by measuring the emitted particle pair correlation function (CF) at small relative momentum. In a basic picture, the CF is the result of the low-energy scattering between the particle pair emitted from the source, and carries the information of the interaction strength, defined by the scattering length $f_{0}$ and effective range $d_{0}$ between the pair, written as ($f_{0}$, $d_{0}$) in short.
Thus, simultaneous inference of the source distribution and the interaction strength becomes feasible for either particle or antiparticle pairs. Based on the Lednický-Lyuboshitz (LL) model, STAR Collaboration has verified that protons and antiprotons exhibit nearly identical interaction properties, and the extracted Gaussian source sizes for $pp$ and $\bar{p}\bar{p}$ pairs were found to be similar [7]. Specifically, within experimental uncertainties, the parameters ($f_{0}$, $d_{0}$) are equivalent for $pp$ and $\bar{p}\bar{p}$ pairs, consistent with CPT symmetry, a fundamental principle in physical laws. The Gaussian source assumption has become widely adopted in correlation function analysis, from which significant results have been obtained in the studies of particle interactions and source distributions [8, 9, 10, 11].
However, by Gaussian parameterization it is assumed that all the participants are totally randomized in space prior to the freeze-out instant, and the initial state as well as the evolution details of the colliding system are lost. This strong assumption is unfavorable in the context of various recent studies. For two examples, i) high energy HIC has been applied to infer the initial shape of the colliding nuclei [12, 13]. ii) The Ohm’s law is suggestively broken because of the strong field and extremely short evolution time [14]. In both cases it is of importance to probe the spatial evolution of the colliding system. Thus, it is surely nontrivial to abandon the Gaussian assumption.
Despite of the difficulty in solving inverse problem, the direct imaging of the source has been achieved without introducing the analytic parameterization [15, 16, 17]. More recently, the Bayessian-based deblurring method, known as Richardson-Lucy (RL) algorithm [18, 19] which was initially developed for optical imaging [20], has gained prominence in nuclear physics [21, 22, 23, 24, 25]. Particularly, Nzabahimana et al. [22] applied the RL algorithm in source imaging in HICs, highlighting the method’s effectiveness and robustness. In this letter, combining the RL deblurring algorithm and the LL model incorporating the pair interaction, we bring the realisation to solve the inverse problem of source imaging from the experimental CFs. The source distribution and ($f_{0}$, $d_{0}$) are determined simultaneously. The similarity of the source spatial distribution of $p$ and $\bar{p}$ emission has been demonstrated.
Defined as a function of the relative momentum $\mathbf{q}=\mathbf{p}_{1}-\mathbf{p}_{2}$ in the particle pair rest frame (PRF), the CF is given experimentally by $C(\mathbf{q})=\frac{P(\mathbf{p}_{1},\mathbf{p}_{2})}{P(\mathbf{p}_{1})P(\mathbf{p}_{2})}$, where $P(\mathbf{p}_{1})$ and $P(\mathbf{p}_{2})$ are the probabilities of detecting particles with momenta $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$, respectively, and $P(\mathbf{p}_{1},\mathbf{p}_{2})$ is the joint probability of detecting both particles simultaneously. For uncorrelated emissions, $C=1$ is expected. Theoretically, the CF can be calculated using the Koonin-Pratt (KP) formula [26, 27]
$C(\mathbf{q})=\int d^{3}r|\Psi_{\mathbf{q}}(\mathbf{r}^{\prime})|^{2}S(\mathbf{r})$ (1)
where $S$ is the source function (SF) related to the relative distance $\mathbf{r}=\mathbf{r}_{1}-\mathbf{r}_{2}$ of the particle-pair in the source rest frame, and $\Psi_{\mathbf{q}}(\mathbf{r}^{\prime})$ is the wave function calculated from the relative momentum $\mathbf{q}$ and distance $\mathbf{r}^{\prime}$ in the PRF. Based on the LL model [28], the interaction strength between particle pair is characterized by ($f_{0}$, $d_{0}$) describing the strong interaction scattering amplitude in an S-wave expansion under the effective range approximation.
Figure 1: Diagram of RL algorithm process.
In optical deblurring, the measured distribution $f(\nu)$ can be related to the true photon distribution $\mathcal{F}(\mu)$ using a conditional probability $P(\nu|\mu)$,
[…]
where the response function $P(\nu|\mu)$ represents the conditional probability that a photon with the true property $\mu$ is measured as $\nu$. Although the true distribution $\mathcal{F}(\mu)$ is unknown, it can be inferred from the measured distribution $f(\nu)$ through iterative regression starting from an arbitrary initial guess. This iterative process is illustrated in Fig. 1. Starting with $t=0$, the arbitrary initial trial $\mathcal{F}^{(0)}(\mu)$ is convolved with $P(\nu|\mu)$ to predict the measurement distribution $f^{P}(\nu)$, which is then compared to the known measurement distribution $f(\nu)$. Based on the difference between them, an amplification factor $A^{(t)}$ is calculated at each $t$th iteration. The factor $A^{(t)}$ is functioning to adjust the values of $\mathcal{F}^{(t)}(\mu)$ for the next iteration. Through the iterative process, the updated distribution $\mathcal{F}^{(t+1)}(\mu)$ progressively converges toward the true distribution $\mathcal{F}(\mu)$. The convergence criteria is invoked if $f^{P}(\nu)$ becomes sufficiently close to $f(\nu)$.
From the mathematical similarity between the KP formula (1) and the optical deblurring equation (2), it is naturally suggested that the RL algorithm can be applied iteratively to solve the SF. By evaluating the converged $\chi^{2}$ value, one can identify the optimal ($f_{0}$, $d_{0}$) for the wave function that best describes the CF, thereby achieving both SF imaging and extraction of strong interaction parameters simultaneously. For the convenience of numerical implementation, the KP formula is discretized by dividing $C$ into $N$ bins and $S$ into $M$ bins ($N\geq M$). The CF $C$ is expressed by $S$ through an $N\times M$ transformation matrix $K$ as $C_{i}=\sum_{j=1}^{M}K_{ij}S_{j}$. This transformation matrix is calculated by
[…]
where $\mathbf{Q}$ denotes the total momentum of the particle pair in the source rest frame. Here, $i$ and $j$ label the bin indexes of $C$ and $S$ , respectively. In one dimensional imaging, the wave function is assumed to be uniform in the polar $\theta$ and azimuthal $\varphi$ directions, leaving only the $r$-direction to consider. Consequently, the integration over $\theta$ and $\varphi$ can be performed using Monte Carlo sampling uniformly within the spherical shell. This approach allows for iterative reconstruction of the source distribution from the discretized experimental measurements of $C$ (see Methods for a detailed description of the iterative process).
Figure 2: (Color Online) Model test procedure. (a) The correlations functions. The blue solid curve is the CF calculated using LL model. The symbols are the sampled CF taking into account the uncertainty to mimic the experimental measurement which is selected as the central CF for imaging. The dashed curves are the output at various iteration steps. The red solid and black dashed thin curves are the final reconstructed CF with two different initial trials of the source function SF1 and SF2. (b) The restored source images. The blue solid curve is the given SF to calculate the CF. The red and black symbols are the restored SF at two initial trials SF1 and SF2. (c) The $\chi^{2}/{\rm NDF}$ as a function of iteration times. (d) The minimum $\chi^{2}/{\rm NDF}$ on the ($f_{0}$, $d_{0}$) plane. The yellow star and red cross denote the input value and the final converged solution.
The imaging of the source and the extraction of interaction parameters ($f_{0}$,$d_{0}$) can now be achieved simultaneously using this method. To test the algorithm, we start with a CF calculated from the given source distribution and known $pp$ interactions with $f_{0}=7.8$ fm and $d_{0}=2.77$ fm for $pp$ interaction. In order to mimic the real measurement, the fluctuation is introduced to each point according to the experimental uncertainty. Fig. 2 (a) shows the theoretic CF (thick solid curve) and the fluctuated CF (open circle). Then the inverse problem is iteratively solved from a initial trial of the SF, which is taken as a Gaussian distribution with $\mu=0$ center location and arbitrary deviation. The reconstructed SF ends close to the known one, as marked SF1 in  2 (b). In order to check the robustness of algorithm against the variation of the initial SF, we also test the second initial trial, where the center of the initial Gaussian is arbitrarily set at $\mu=3$ fm instead of zero peaking. With the iteration going on, the output SF gradually approaches to the known one, as illustrated by the dashed curves in Fig.  2 (b). The marked SF2 sits on top of SF1, evidencing the robustness of the iterative procedure. Fig. 2 (c) presents the closeness quantity $\chi^{2}$ over iteration times in the imaging process of the central CF. Here $\chi^{2}$ is defined as $\chi^{2}=\sum_{i}^{N}\left(\frac{C_{i}^{P}-C_{i}}{\sigma_{i}}\right)^{2}$, where $C_{i}^{P}$ represents the value of the $i^{\rm th}$ point. During the procedure, ($f_{0}$,$d_{0}$) parameters are optimized by minimum $\chi^{2}$ and the closeness between the output and the known CF is checked in every iteration till the convergence condition is reached, namely the change of $\chi^{2}$ between two successive iterations is less than $1\times 10^{-4}$ , which functions as a threshold to avoid noise amplification. Fig. 2 (d) shows the minimum $\chi^{2}$ distribution on the ($f_{0}$,$d_{0}$) plane. The optimized ($f_{0}$,$d_{0}$), marked by red cross, converge to the input one (yellow star). The consistency between the input values and the reconstructed ones in panel (a), (b) and (d) demonstrates the validity of the deblurring algorithm.
Now we apply the method to the $pp$ and $\bar{p}\bar{p}$ CFs in Au+Au at center-of-mass energy of $200~{}\rm{GeV}$ per nucleon pair [29], taken by STAR collaboration [7, 30] at the Relativistic Heavy Ion Collider (RHIC) [31, 32]. The experimental CFs are inclusive and contain the contributions from the residual CFs of $p\Lambda$ and $\Lambda\Lambda$. The residual contribution, which shall be subtracted, is derived by the difference between the two fitting curves with and without the feedback corrections shown in [29]. The pure $pp$ and $\bar{p}\bar{p}$ CFs are presented in Fig. 3 (a) by the symbols, where the uncertainties are taken same as in the experiment for each data point.
In the calculation of the Kernel $K_{ij}$, one needs the momentum distribution of $p$ and $\bar{p}$. The transverse momentum $p_{\rm t}$ distributions are taken from the midrapidity range in the same colliding system [33] measured by PHENIX experiment [34, 35]. The longitudinal rapidity $y_{\rm l}$ as well as the azimuth in transverse direction are uniformly distributed. The cuts $0.4<p_{\rm t}<2.5\rm{GeV}/c$ and $|y_{\rm l}|<0.7$ are taken as same as STAR data where the CFs are measured. As well noticed, the LL model provides relatively weak constraints on $d_{0}$ when analyzing experimental data. So the widely accepted theoretical and experimental approximation $d_{0}=2.8\rm~{}fm$ is fixed here for both $pp$ and $\bar{p}\bar{p}$ pairs. The convergence criterion is set to $1\times 10^{-2}$. The corresponding optimal values of $f_{0}$ obtained are as follows:
$f_{0(pp)}=(7.7\pm 0.2)\text{fm},\quad f_{0(\bar{p}\bar{p})}=(7.9\pm 0.3)\text{fm}.$ (4)
The $f_{0}$ values for $pp$ and $\bar{p}\bar{p}$ pairs are identical and consistent with both the theoretical calculations and the experimental results from STAR collaboration [7].
Figure 3: (Color Online) (a) The reconstructed CFs of $pp$ (red solid curve) and $\bar{p}\bar{p}$ (blue dashed curve) pairs in comparison to the experimental CFs (red dots for $pp$ and blue circles for $\bar{p}\bar{p}$ pairs). (b) Reconstructed SFs for $p$ (red dots) and $\bar{p}$ (blue circles) in Au+Au at center of mass energy 200 GeV. The symbols are the results of direct imaging without Gaussian assumption. The SFs with Gaussian shape obtained in [7] are shown by the curves for reference.
The converged CFs of $pp$ and $\bar{p}\bar{p}$ reproduce fairly the experimental data, as demonstrated by the solid and dashed curves in Fig. 3 (a), respectively. The results of the restored source imaging are plotted in panel (b) by red dots for $p$ and blue circles for $\bar{p}$. In order to estimate the uncertainty of each imaging point, numerous pseudo-CFs for $pp$ and $\bar{p}\bar{p}$ are sampled, smearing the content in each individual $q$ bin by Gaussian fluctuation, for which the center (standard deviation) is taken from the experimental value (uncertainty). Then the standard deviation of the reconstructed SFs are calculated and plotted on each data point bin-by-bin as the error bar on the SF.
The direct imaging result in Fig. 3 (b) has important implications. First, Within the uncertainty, protons and antiprotons share the identical source distribution at the instant of freeze-out. In addition to the momentum distributions done in [29], the direct imaging result here completes the last piece of evidence of the matter-antimatter symmetry in coordinate space as they are produced and evolve in high energy HICs. This evidence is obtained without imposing assumptions to the shape of the source. Second, the profile of the reconstructed SFs deviates from the Gaussian distribution for both protons and antiprotons. Quantitative calculation yields the probability for the SF to be Gaussian is 00 and $2.2\times 10^{-14}$ for $p$ and $\bar{p}$, respectively. For a clear comparison, the Gaussian distribution obtained in [29] are shown by the curves in panel (b), respectively. It is evident that the emission source are more populated at smaller radius ($r<5$ fm) but less distributed at larger radius ($105<r<$ 10 fm), as zoomed in in the inset. It is consistent with the picture of an extremely fast collision. Namely, protons and antiprotons, formed by the coalescence of quarks and antiquarks created inside the overlapping region, are not fully randomized in coordinate space prior to the freeze-out instant, otherwise they are expected to extend to the outer space in Gaussian distribution. It offers a support to the recent idea of probing the initial shape of the nucleus via relativistic HIC [36, 37, 38, 39, 13, 40], since the collision is so rapid that the spatial information of the colliding nuclei is partially preserved.
The impact of the bin-by-bin variation of the SF on the CF has been analyzed. It is observed that variations in the first few bins of the source function (except for the first bin which is below the low limit cut in LL model calculation) significantly impact the CF, whereas the variations at larger $r$ values are less influential. Inversely, the experimental uncertainty of the CFs near the peak ($q\approx 40~{}\rm MeV$) brings considerable variation to the SF. In order to achieve more accurate imaging, it is demanded to further reduce the uncertainty of the CF, particularly at small $q$. If given high-accuracy experimental CFs, the direct imaging technique developed here offers a novel opportunity to probe the proton density profile of the exotic bubble nucleus, which has been recently proposed [41] but yet in debate [42], in the relativistic collisions induced by those nuclei on future radioactive ion beam facilities. Nevertheless, due to the non-negativity constraint required by the Bayesian-based deblurring algorithm, the method is limited by the choice of basis functions for representing the source. Future investigations are required to explore new inversion techniques such as neural networks, and to adopt mathematically more suitable source bases, as well as to optimize the selection of the convergence criteria, which will all advance the development and application of imaging techniques based on correlation function method.
To summarize, a direct imaging method on femtometer scale has been realized by combining the Richardson-Lucy deblurring algorithm and the intensity interferometry in ultra relativistic heavy ion collisions. This method allows for a more accurate extraction of source distributions and final state interaction between particle pair based on the Lednický-Lyuboshitz model analysis. Applying this imaging method to $pp$ and $\bar{p}\bar{p}$ correlation functions in Au+Au at center-of-mass energy 200 GeV measured by the STAR collaboration, we observe the same freeze-out source functions for $p$ and $\bar{p}$ emissions with distinct non-Gaussian characteristics. These findings confirm the matter-antimatter symmetry in coordinate space once produced and suggest the ultrafast feature of the evolution of the collision, which is hardly accessible through other experimental means alone. The new method can be extended to three dimensional imaging. It represents a significant advancement of our ability to image the femtoscopic heavy ion collisions and to explore matter-antimatter physics in terrestrial laboratories.
Acknowledgements  This work is supported by the National Natural Science Foundation of China under Grant Nos.12335008 and 12205160. The authors acknowledge the support of the Center for High Performance Computing and Initiative Scientific Research Program in Tsinghua University. Sincere gratitude is extended to Nu Xu and Pengfei Zhuang for their valuable suggestions and insightful discussions.
Author contributions  JHX, code developing and calculation. ZQ, code reviewing. YJW, analysis and result checking. ZGX, project leading and supervising. All authors contribute to discussions, draft writing and editing.